OpenStudy (anonymous):

find the derivative of e^(1/x) + 1/(ex^2) don't quite know how to solve this :/

zepdrix (zepdrix):

Hmm this one isn't too bad :) Ooo a dolphin!

zepdrix (zepdrix):

Is this what the problem is supposed to look like?\[\Large\rm e^{1/x}+\frac{1}{e^{x^2}}\]Or this?\[\Large\rm e^{1/x}+\frac{1}{ex^2}\]

zepdrix (zepdrix):

Or neither? :U

OpenStudy (anonymous):

uh, its the first one @zepdrix

zepdrix (zepdrix):

Let's apply our exponent rule to write the expression like this:\[\Large\rm e^{1/x}+e^{-x^2}\]

zepdrix (zepdrix):

Do you remember your derivative of e^x?

OpenStudy (anonymous):

yes its e^x

zepdrix (zepdrix):

mmm ok good. So e^x gives us e^x. Gives us the same thing back. Same idea for more complex exponentials like this one, we simply have to apply the chain rule on top of that.

zepdrix (zepdrix):

\[\Large\rm \frac{d}{dx}e^{1/x}=e^{1/x}\left(\frac{d}{dx}\frac{1}{x}\right)\]The derivative of the exponential gave us the same thing back. Chain rule tells us to multiply by the derivative of the inner function.

zepdrix (zepdrix):

If you forgot how to differentiate \(\Large\rm \frac{1}{x}\), you can rewrite it as \(\Large\rm x^{-1}\) and apply your power rule.

OpenStudy (anonymous):

okay wait so can you write it out so i can see exactly what you mean by that

zepdrix (zepdrix):

I guess you need to remember this with your exponent rules:\[\Large\rm x^{-a}=\frac{1}{x^a}\]The negative tells us to flip it into the denominator. In our problem, we're doing that in reverse.\[\Large\rm \frac{1}{x^1}=x^{-1}\]

OpenStudy (anonymous):

yes i understand that part but are u saying that the e^(1/x) stays the same?

zepdrix (zepdrix):

yes, that portion stays the same. Only need to apply chain rule.

OpenStudy (anonymous):

okay so then the answer would be e^(1/x)-2ex^(-3) ?

zepdrix (zepdrix):